Systems of equations: solve using elimination method
X+2y-z = -6
3x-y+z = 8
4x-y-2z = 0
Answer #1

Xtay-2 =-6 - o 3x -y+z= 8 - w 4x -Y-22 =o - from eq 0 4 ③ Xtay-Z=-6 3X - Y + 2 = 8 - 4X+Y=2 - on butting the value of x in eq